OpenStudy (anonymous):

1 more question and im done :)...-m2k+(-km2)..the 2 mean squared

OpenStudy (anonymous):

help!! sigh sigh

OpenStudy (anonymous):

wow this questions hard?no one can help me........

myininaya (myininaya):

\[-m^2k+(-km^2)\] is that right?

OpenStudy (anonymous):

yea sry

myininaya (myininaya):

well we could write it as \[-m^2k-m^2k\]

myininaya (myininaya):

there terms are the same 1+1=2 -1-1=-2 so we have \[-2m^2k\]

OpenStudy (anonymous):

ty sister u are cool!! hope ull success in ur life :)